YEAR 1 (R-Billion) YEAR 2 (R-Billion)

Investment 200 220



Saving 180 190

Export 100 110

Imports 120 140

Government Expenditure 150 160

Taxation 150 160

Equilibrium National Income 1 800 2 000



(a) Calculate the value of the multiplier for this economy.

(b) Should the full employment level of income be R-B2 255, by how much should government change its spending to reach this level of income during the next year given the multiplier has not changed.

(c) Evaluate whether or not this policy approach is effective in real life in achieving the desired level of GDP.
1
Expert's answer
2021-07-26T17:28:01-0400

Solution:

a.). Multiplier:

Multiplier = "\frac{1}{1 - MPC}"


MPC = "\frac{\triangle C}{\triangle Y}"


First determine consumption:

Consumption = Income – savings

Consumption Year 1 = 1800 – 180 = 1,620

Consumption Year 2 = 2000 – 190 = 1,810

Change in consumption = 1810 – 1620 = 190

Change in income = 2000 – 1800 = 200

MPC ="\frac{190}{200} = 0.95"


The multiplier for this economy = 0.95


b.). Let the new income level be 2, 255 and we use Year 2 data:

Y = C + I + G + X-M

C = 1810 + 0.95(Y – T) = 1810 + 0.95(2255 – 160) = 3800.25

2255 = 3800.25 + 220 + G + 110 – 140

2255 = 3990.25 + G

G = 3990.25 – 2255 = 1735.25

Previous Government spending = 160

Change = 1735.25 – 160 = 1575.25

Therefore, the government should increase its spending by 1,575.25

 

c.). This policy approach is effective in real life in achieving the desired level of GDP. This is because the government can use this policy by shifting the aggregate demand to the right to come out of recession to maintain a high rate of economic growth and stabilize prices and wages.